OpenStudy (anonymous):

y=cos^7(sin3x) I think i figured out the right answer: -21cos^6x(sin3x)sin(sin3x)cos(3x) I'm just trying to input it into my online homework and it keeps saying its incorrect, what am i doing wrong!?!?!

OpenStudy (jhannybean):

What exactly are you trying to do with this equation?

OpenStudy (anonymous):

\[y=\cos^7(\sin3x)\]\[y=[\cos~(\sin3x)]^7\]taking the derivative, because that is what you seem to try to do. \[y' =7[\cos~(\sin3x)]^6\times (-\sin~(\sin3x)~)\times (~\cos(3x)~)\times 3\]\[y' =-63[\cos^6(\sin3x)](\sin~(\sin3x)~)\times (~\cos(3x)~)\]you forgot the chain rule for 3x.

OpenStudy (anonymous):

there is a chain rule 3 times. and I hope you figure out what you left out....

OpenStudy (anonymous):

thanks!